# Data-Centric Foundation Models in Computational Healthcare: A Survey

### v1

@@ 7. Data Privacy
Safeguarding healthcare data privacy is crucial given their critical roles in clinical research, diagnosis, treatment, and disease prevention (Gkoulalas-Divanis et al., 2014). Laws and regulations have been enacted to ensure the protection of these highly sensitive personal records (Clayton et al., 2023; Malin et al., 2010), fostering the trust between patients and medical institutions.
@@
Traditionally, healthcare data is well protected to preserve privacy by strategies such as cryptographic methods and access control mechanisms (Gertz and Jajodia, 2007; Aggarwal and Yu, 2008). Nevertheless, these approaches encounter limitations in providing absolute patient anonymity, whereby the private and confidential information of patients remains undisclosed, particularly when disseminating patient data (Fung et al., 2010). The challenge arises due to the necessity of distributing data to a broad and potentially unidentified group of recipients. To solve this, healthcare data is typically altered to remove private information before being used (de Wolf, 2012; Willenborg and De Waal, 2012; Kosorok and Laber, 2019). Perturbation methods are used to add controlled noise or randomness to data to protect privacy while allowing data analysis (Turgay et al., 2023). These methods are leveraged at a reduced level of intensity to mitigate the potential deterioration in the integrity of subsequent secondary data analyses (Reiter, 2012). However, with the emergence of accessible digital data and powerful computational resources, attackers can now obtain huge amounts of information to link released records to external data files (Reiter, 2019). This critical concern leads the research to generate synthetic data that contain no actual values of individual information (Reiter, 2023).
@@
The advent of data-centric FMs introduces both opportunities and challenges in healthcare data privacy. On one hand, FMs offer improved capabilities for data generation, providing unidentified synthetic data (Tang et al., 2023a; Shibata et al., 2023). On the other hand, FMs tend to memorize their training data due to their massive parameters, resulting in the generation of similar outputs that potentially contain private patient information (Zhou et al., 2023b; Carlini et al., 2021).
@@
FMs offer a practical opportunity for addressing healthcare data privacy concerns. The growing use of synthetic data represents a solution to protect private information that is in line with the key promise of FMs (Hu et al., 2023a). Deep generative models can now produce high-quality synthetic medical data with similar characteristics to the original data but without identical information (Naresh et al., 2023), thus preserving the functionality while avoiding risks of privacy leakage. For instance, recent advances in FM-based data generation methods have enabled the anonymization of medical images and texts, supporting data privacy while maintaining a wide utility (Dong et al., 2022b). FMs also enable the generation of differentially private data for model training (Sei et al., 2024), which ensures the effectiveness of model training, while the generated synthetic data preserves data privacy. Shibata et al. (Shibata et al., 2023) propose a large-scale diffusion model capable of unconditionally generating high-resolution volumetric medical images, which removes critical information while remaining trainable features, preserving privacy when sharing data in various domains beyond healthcare. Similarly, the use of synthetic imaging data has shown its promise in protecting patient records while retaining high-level segmentation performance in a multi-center evaluation (Chang et al., 2023c).
@@
Although FMs shed light on addressing medical data privacy issues, we have recognized that FMs themselves have brought up emerging privacy protection challenges. Numerous studies have extensively investigated FMs of their propensity to memorize training data across different domains, a phenomenon referred to as memorization (Carlini et al., 2019; Thakkar et al., 2021; Zhang et al., 2021b; Carlini et al., 2022; Kandpal et al., 2022; Lee et al., 2021). FMs have been trained on private datasets (Carlini et al., 2023). This characteristic of FMs makes them prone to generating text sequences or other information directly from their training data (Carlini et al., 2021), which may include sensitive patient information. Zhou et al. (Zhou et al., 2023b) highlight the possibility of retrieving specific training samples by querying the massive FMs, leading to private information leakage. Even worse, due to the billion parameters of FMs, they are more likely to retain a greater amount of private information, making the larger model more vulnerable than smaller models. Recently, more researchers have paid attention to FM privacy issues in healthcare. The memorization phenomenon poses a risk of capturing and potentially revealing highly sensitive information, such as protected health information (Lehman et al., 2021; Pan et al., 2020). Sallam et al. (Sallam et al., 2023) highlight data privacy concerns associated with ChatGPT-generated content in healthcare education, emphasizing the need for strict regulation and full protection of patient information. To mitigate the risks of privacy breaches, it is essential to consider implementing safeguards when utilizing FMs in healthcare. These safeguards should encompass various stages such as data processing, model training, model inference, and system deployment.
@@ 8.1. Benchmarking
Researchers have established benchmarks for model evaluation in the healthcare field, inspired by the progress of AI research in vision and language domains (Mincu and Roy, 2022; Johnson et al., 2016; Gu et al., 2021). A remarkable feature is that these benchmarks often build on sample diversity and large quantity scale to facilitate FM evaluation. For example, The Biomedical Language Understanding Evaluation benchmark (BLUE) (Peng et al., 2019) covers five language tasks with ten biomedical and clinical text datasets, mixing PubMed and MIMIC-III (Johnson et al., 2016) based applications. The Biomedical Language Understanding & Reasoning Benchmark (BLURB) (Gu et al., 2021) focuses on biomedical datasets and improves BLUE by including question-answering tasks which are frequently employed to evaluate large language models (LLMs). Table 1 displays a representative list of public medical databases, benchmarks, and datasets that can facilitate FM training, inference, and evaluation.
@@
Database/Benchmark/Dataset Description Task type MIMIC-III (Johnson et al., 2016) Critical care data for over 40,000 patients Text BLUE (Peng et al., 2019) 5 language tasks with 10 biomedical and clinical text datasets Text MedMentions (Mohan and Li, 2019) 4,392 papers annotated by experts with mentions of UMLS entities Text webMedQA (He et al., 2019) 63,284 real-world Chinese medical questions with 300K answers Text PubMedQA (Jin et al., 2019) 1K expert-annotated, 61.2K unlabeled, and 211.3K artificially generated biomedical QA instances Text BLURB (Gu et al., 2021) 13 biomedical NLP datasets in 6 tasks Text CBLUE (Zhang et al., 2021a) A Chinese medical NLP benchmark with 18 datasets Text MedQA-USMLE (Jin et al., 2021) 61,097 multiple choice questions based on USMLE in three languages Text MedMCQA (Pal et al., 2022) 194K multiple-choice questions covering 2.4K healthcare topics Text MultiMedQA (Singhal et al., 2022) 6 existing and 1 online-collected medical QA dataset Text ClinicalTrials.gov [^3] https://clinicaltrials.gov/ An online database of clinical research studies, including clinical trials and observational studies Text Medical Meadow (Han et al., 2023) 16M medical QA pairs collected from 9 sources Text Huatuo-26M (Li et al., 2023d) 26M Chinese medical QA pairs Text ChestXray-NIHCC (Wang et al., 2017) 100K radiographs with labels from more than 30,000 patients Imaging DeepLesion (Yan et al., 2018) 32K CT scans with annotations and semantic labels from radiological reports Imaging Kather Colon Dataset (Kather et al., 2018) 100K histological images of human colorectal cancer and healthy tissue Imaging CheXpert (Irvin et al., 2019) 224,316 chest radiographs of 65,240 patients Imaging EchoNet-Dynamic (Ouyang et al., 2020) 10,030 expert-annotated echocardiogram videos Imaging ISIC [^4] https://www.isic-archive.com/ An archive containing 23K skin lesion images with labels Imaging Med-MNIST v2 (Yang et al., 2023a) 12 2D and 6 3D datasets for biomedical image classification Imaging AbdomenAtlas-8K (Qu et al., 2023) 8,448 CT volumes with per-voxel annotated eight abdominal organs Imaging Virchow (Vorontsov et al., 2023) 1.5M pathological scans from 120K patients Imaging RETFound (Zhou et al., 2023a) Unannotated retinal images, containing 904,170 CFPs and 736,442 OCT scans Imaging dbSNP (Sherry et al., 2001) A collection of human single nucleotide variations, microsatellites, and small-scale insertions and deletions Genomics ENCODE (Consortium et al., 2012) A platform of genomics data and encyclopedia with integrative-level and ground-level annotations Genomics 1000 Genomes Project (Consortium et al., 2015) A comprehensive catalog of human genetic variations Genomics ChEMBL (Gaulton et al., 2012) 20M bioactivity measurements for 2.4M distinct compounds and 15K protein targets Drug DrugBank (Wishart et al., 2018) A web-enabled structured database of molecular information about drugs Drug PubChem (Kim et al., 2023) A collection of 900+ sources of chemical information data Drug TCGA [^5] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A landmark cancer genomics program, molecularly characterized over 20,000 primary cancer and matched normal samples spanning 33 cancer types Multi-modal MIMIC-CXR (Johnson et al., 2019) 227,835 chest imaging studies with free-text reports for 65,379 patients Multi-modal MIMIC-IV (Johnson et al., 2020) Clinical information for hospital stays of over 60,000 patients Multi-modal SwissProtCLAP (Liu et al., 2023g) 441K text-protein sequence pairs Multi-modal PMC-VQA (Zhang et al., 2023d) 227K VQA pairs of 149K images of various modalities or diseases Multi-modal MedMD (Wu et al., 2023e) 15.5M 2D scans and 180k 3D radiology scans with textual descriptions Multi-modal PathCap (Sun et al., 2023) 142K pathology image-caption pairs from various sources Multi-modal PathInstruct (Sun et al., 2023) 180K samples of LLM-generated instruction-following data Multi-modal OpenPath (Huang et al., 2023a) 208,414 pathology images paired with natural language descriptions Multi-modal Quilt-1M (Ikezogwo et al., 2023) 1M image-text pairs for histopathology Multi-modal PMC-OA (Lin et al., 2023b) 1.6M fine-grained biomedical image-text pairs Multi-modal
@@
Benchmarks also cover different evaluation metrics for a wide range of tasks. Conventional evaluation metrics include accuracy, area under the curve (AUC), and mean average precision (mAP) for classification, intersection over union (IoU) for image segmentation, and bilingual evaluation understudy (BLEU) and BERTScore (Zhang et al., 2019) for text-related tasks, which mostly emphasize precise model predictions. However, FM evaluation metrics largely go beyond mere accuracy. Recent work has queried whether benchmark evaluation metrics align with human values (Wornow et al., 2023; He et al., 2023b). The lack of such alignment could pose misleading estimates of FM capabilities leading to harmful results. In fact, LLM alignment evaluation from perspectives of reliability, safety, fairness, explainability, and robustness underscores the importance of LLM trustworthiness (Liu et al., 2023d). Yet we lack substantial research benchmarks for FM trustworthiness evaluation in the healthcare field.
@@ 8.3. Automated Evaluation
Leveraging powerful FMs for automated evaluation is a promising research direction. A high-performing and well-aligned FM can be employed as a benchmark or an evaluation tool to automatically assess other FMs, eliminating the need of static benchmarks or human experts. For instance, Chen et al. (Chen et al., 2023c) explore reference-free evaluation methods that prompt ChatGPT to score the quality of model-generated texts without a pre-defined ground truth. Chiang et al. (Chiang and Lee, 2023) show that LLM evaluation can be stable over prompt formatting and consistent with human experts. Jain et al. (Jain et al., 2023) propose self-supervised evaluation strategies to assess LLM properties without benchmarks or human annotations. Ye et al. propose FLASK (Ye et al., 2023), a fine-grained LLM evaluation protocol that decomposes coarse-level scoring to instance-wise skill assessment. FLASK prompts GPT-4 to assign a score with a rationale for each alignment skill of the target FM-generated text instance. From a data-centric perspective, the benchmark FM with extensive knowledge and human alignment embedded in its massive pre-training data can evaluate the performance of other FMs.
@@ 9.3. Bias
From a model perspective, the inherent bias in FM architectures should be recognized by healthcare stakeholders and developers. Currently, FM architectures and training patterns lack the defense and detection capability against adversarial manipulation. FMs could propagate all knowledge learned from training data without any regulation of biased knowledge detection. This vulnerability can potentially hurt healthcare stakeholders by generating false or harmful information. Meanwhile, adversarial attack training is promising to enhance the harmful information defense and detection capability (Ferrara, 2023). Furthermore, the bias of human-AI alignment remains in FM development, where matching human value and model output will be a long-term challenge. To mitigate this bias, strong guidance from human experts in dataset curation and FM outcome evaluation is necessary in the healthcare domain. Human experts can provide reliable feedback to FM developers for identifying and addressing bias-related issues (Ferrara, 2023).

### v2

@@ 7. Data Privacy
Safeguarding healthcare data privacy is crucial given their critical roles in clinical research, diagnosis, treatment, and disease prevention (Gkoulalas-Divanis et al., 2014). Laws and regulations have been enacted to safeguard these highly sensitive personal records (Clayton et al., 2023; Malin et al., 2010), which helps to build trust between patients and medical institutions.
@@
Traditionally, healthcare data is well protected to preserve privacy through techniques like encryption and access control mechanisms (Gertz and Jajodia, 2007; Aggarwal and Yu, 2008). Nevertheless, these approaches encounter limitations in providing absolute patient anonymity, whereby private and confidential information of patients remains undisclosed, particularly when disseminating patient data (Fung et al., 2010). The challenge arises due to the necessity of distributing data to a broad and potentially unidentified group of recipients. To solve this, healthcare data is typically altered to remove private information before being used (de Wolf, 2012; Willenborg and De Waal, 2012; Kosorok and Laber, 2019). Perturbation methods involve adding controlled noise or randomness to data to protect privacy while allowing data analysis (Turgay et al., 2023). These methods are leveraged at a reduced level of intensity to mitigate the potential deterioration in the integrity of subsequent secondary data analyses (Reiter, 2012). However, with the emergence of accessible digital data and powerful computational resources, attackers can now obtain huge amounts of information to link released records to external data files (Reiter, 2019). This critical concern leads the research to generate synthetic data that contain no actual values of individual information (Reiter, 2023).
@@
The advent of data-centric FMs introduces both opportunities and challenges in healthcare data privacy. On the one hand, FMs offer improved capabilities for data generation, providing unidentified synthetic data (Tang et al., 2023a; Shibata et al., 2023). On the other hand, FMs tend to memorize their training data due to their massive parameters, resulting in the generation of similar outputs that potentially contain private patient information (Zhou et al., 2023b; Carlini et al., 2021).
@@
FMs offer a practical opportunity for addressing healthcare data privacy concerns. The growing use of synthetic data represents a solution to protect private information that is in line with the key promise of FMs (Hu et al., 2023a). Deep generative models can now produce high-quality synthetic medical data with similar characteristics to the original data but without identical information (Naresh et al., 2023), thus preserving functionality while avoiding privacy leakage. For instance, recent advances in FM-based data generation methods have enabled the anonymization of medical images and texts, supporting data privacy while maintaining a wide utility (Dong et al., 2022b). FMs also enable the generation of differentially private data for model training (Sei et al., 2024), which ensures the effectiveness of model training, while the generated synthetic data preserves data privacy. Shibata et al. (Shibata et al., 2023) propose a large-scale diffusion model capable of unconditionally generating high-resolution volumetric medical images, which removes critical information while remaining trainable features, preserving privacy when sharing data in various domains beyond healthcare. Similarly, the use of synthetic imaging data has shown its promise in protecting patient records while retaining high-level segmentation performance in a multi-center evaluation (Chang et al., 2023c).
@@
Although FMs shed light on addressing medical data privacy issues, we have recognized that FMs themselves have brought up emerging privacy protection challenges. Numerous studies have extensively investigated FMs of their propensity to memorize training data across different domains, a phenomenon referred to as memorization (Carlini et al., 2019, 2022; Kandpal et al., 2022; Lee et al., 2021). FMs have been trained on private datasets (Carlini et al., 2023). This characteristic of FMs makes them prone to generating text sequences or other information directly from their training data (Carlini et al., 2021), which may include sensitive patient information. Zhou et al. (Zhou et al., 2023b) highlight the possibility of retrieving specific training samples by querying the massive FMs, leading to private information leakage. Even worse, due to the billion parameters of FMs, they are more likely to retain a greater amount of private information, making the larger model more vulnerable than smaller models. Recently, more researchers have paid attention to FM privacy issues in healthcare. The memorization phenomenon poses a risk of capturing and potentially revealing highly sensitive information, such as protected health information (Lehman et al., 2021; Pan et al., 2020). Sallam et al. (Sallam et al., 2023) highlight data privacy concerns associated with ChatGPT-generated content in healthcare education, emphasizing the need for strict regulation and full protection of patient information. To mitigate the risks of privacy breaches, it is essential to consider implementing safeguards when utilizing FMs in healthcare. These safeguards should encompass various stages such as data processing, model training, model inference, and system deployment.
@@ 8.1. Benchmarking
Researchers have established benchmarks for model evaluation in the healthcare field, inspired by the progress of AI research in vision and language domains (Mincu and Roy, 2022; Johnson et al., 2016; Gu et al., 2021). A remarkable feature is that these benchmarks often build on sample diversity and large quantity scale to facilitate FM evaluation. For example, The Biomedical Language Understanding Evaluation benchmark (BLUE) (Peng et al., 2019) covers five language tasks with ten biomedical and clinical text datasets, mixing PubMed and MIMIC-III (Johnson et al., 2016) based applications. The Biomedical Language Understanding & Reasoning Benchmark (BLURB) (Gu et al., 2021) focuses on biomedical datasets and improves BLUE by including question-answering tasks which are frequently employed to evaluate large language models (LLMs). Medbench (Liu et al., 2024a) provides over 300K Chinese questions covering 43 clinical specialties, an automatic evaluation system, and dynamic evaluation mechanisms for unbiased and reliable assessment of Chinese medical LLMs. Table 1 displays a representative list of public medical databases, benchmarks, and datasets that can facilitate FM training, inference, and evaluation.
@@
Database/Benchmark/Dataset Description Task type ClinicalTrials.gov [^3] https://clinicaltrials.gov/ An online database of clinical research studies, including clinical trials and observational studies Text MIMIC-III (Johnson et al., 2016) Critical care data for over 40,000 patients Text BLUE (Peng et al., 2019) 5 language tasks with 10 biomedical and clinical text datasets Text MedMentions (Mohan and Li, 2019) 4,392 papers annotated by experts with mentions of UMLS entities Text webMedQA (He et al., 2019) 63,284 real-world Chinese medical questions with 300K answers Text PubMedQA (Jin et al., 2019) 1K expert-annotated, 61.2K unlabeled, and 211.3K artificially generated biomedical QA instances Text BLURB (Gu et al., 2021) 13 biomedical NLP datasets in 6 tasks Text CBLUE (Zhang et al., 2021) A Chinese medical NLP benchmark with 18 datasets Text MedQA-USMLE (Jin et al., 2021) 61,097 multiple choice questions based on USMLE in three languages Text MedMCQA (Pal et al., 2022) 194K multiple-choice questions covering 2.4K healthcare topics Text MultiMedQA (Singhal et al., 2022) 6 existing and 1 online-collected medical QA dataset Text Medical Meadow (Han et al., 2023) 16M medical QA pairs collected from 9 sources Text Huatuo-26M (Li et al., 2023d) 26M Chinese medical QA pairs Text GAP-Replay (Chen et al., 2023a) 48.1B tokens from 4 medical corpora including guidelines, abstracts, papers, and replay Text BiMed1.3M (Pieri et al., 2024) An English and Arabic bilingual dataset of 1.3M medical QA and chat samples Text MMedC (Qiu et al., 2024) A multilingual medical corpus containing over 25.5B tokens Text MedBench (Liu et al., 2024a) 300,901 Chinese questions covering 43 clinical specialties, combined with an automatic evaluation system Text ISIC [^4] https://www.isic-archive.com/ An archive containing 23K skin lesion images with labels Imaging ChestXray-NIHCC (Wang et al., 2017) 100K radiographs with labels from more than 30,000 patients Imaging DeepLesion (Yan et al., 2018) 32K CT scans with annotations and semantic labels from radiological reports Imaging Kather Colon Dataset (Kather et al., 2018) 100K histological images of human colorectal cancer and healthy tissue Imaging CheXpert (Irvin et al., 2019) 224,316 chest radiographs of 65,240 patients Imaging EchoNet-Dynamic (Ouyang et al., 2020) 10,030 expert-annotated echocardiogram videos Imaging Med-MNIST v2 (Yang et al., 2023a) 12 2D and 6 3D datasets for biomedical image classification Imaging AbdomenAtlas-8K (Qu et al., 2023) 8,448 CT volumes with per-voxel annotated eight abdominal organs Imaging Virchow (Vorontsov et al., 2023) 1.5M pathological scans from 120K patients Imaging AbdomenAtlas-8K 8,448 CT volumes with per-voxel annotated eight abdominal organs Imaging RETFound (Zhou et al., 2023a) Unannotated retinal images, containing 904,170 CFPs and 736,442 OCT scans Imaging dbSNP (Sherry et al., 2001) A collection of human single nucleotide variations, microsatellites, and small-scale insertions and deletions Genomics ENCODE (Consortium et al., 2012) A platform of genomics data and encyclopedia with integrative-level and ground-level annotations Genomics 1000 Genomes Project (Consortium et al., 2015) A comprehensive catalog of human genetic variations Genomics ChEMBL (Gaulton et al., 2012) 20M bioactivity measurements for 2.4M distinct compounds and 15K protein targets Drug DrugBank (Wishart et al., 2018) A web-enabled structured database of molecular information about drugs Drug PubChem (Kim et al., 2023) A collection of 900+ sources of chemical information data Drug DrugChat (Liang et al., 2023) 143,517 question-answer pairs covering 10,834 drug compounds, collected from PubChem and ChEMBL Drug TCGA [^5] https://www.cancer.gov/ccg/research/genome-sequencing/tcga A landmark cancer genomics program, molecularly characterized over 20,000 primary cancer and matched normal samples spanning 33 cancer types Multi-modal MIMIC-CXR (Johnson et al., 2019) 227,835 chest imaging studies with free-text reports for 65,379 patients Multi-modal MIMIC-IV (Johnson et al., 2020) Clinical information for hospital stays of over 60,000 patients Multi-modal SwissProtCLAP (Liu et al., 2023h) 441K text-protein sequence pairs Multi-modal PMC-VQA (Zhang et al., 2023d) 227K VQA pairs of 149K images of various modalities or diseases Multi-modal MedMD (Wu et al., 2023e) 15.5M 2D scans and 180k 3D radiology scans with textual descriptions Multi-modal PathCap (Sun et al., 2023) 142K pathology image-caption pairs from various sources Multi-modal PMC-OA (Lin et al., 2023b) 1.6M fine-grained biomedical image-text pairs Multi-modal PathInstruct (Sun et al., 2023) 180K samples of LLM-generated instruction-following data Multi-modal Quilt-1M (Ikezogwo et al., 2023) 1M image-text pairs for histopathology Multi-modal OpenPath (Huang et al., 2023a) 208,414 pathology images paired with natural language descriptions Multi-modal Chi-Med-VL (Liu et al., 2023d) 580,014 image-text pairs and 469,441 question-answer pairs for general healthcare in Chinese Multi-modal SAT-DS (Zhao et al., 2023) 11,462 scans with 142,254 segmentation annotations spanning 8 human body regions Multi-modal
@@
Benchmarks also cover different evaluation metrics for a wide range of tasks. Conventional evaluation metrics include accuracy, area under the curve (AUC), and mean average precision (mAP) for classification, intersection over union (IoU) for image segmentation, and bilingual evaluation understudy (BLEU) and BERTScore (Zhang et al., 2019) for text-related tasks, which mostly emphasize precise model predictions. However, FM evaluation metrics largely go beyond mere accuracy. Recent work has queried whether benchmark evaluation metrics align with human values (Wornow et al., 2023; He et al., 2023b). The lack of such alignment could pose misleading estimates of FM capabilities leading to harmful results. In fact, LLM alignment evaluation from perspectives of reliability, safety, fairness, explainability, and robustness underscores the importance of LLM trustworthiness (Liu et al., 2023e). Yet we lack substantial research benchmarks for FM trustworthiness evaluation in the healthcare field.
@@ 8.3. Automated Evaluation
Leveraging powerful FMs for automated evaluation is a promising research direction. A high-performing and well-aligned FM can be employed as a benchmark or an evaluation tool to automatically assess other FMs, eliminating the need of static benchmarks or human experts. For instance, Chen et al. (Chen et al., 2023e) explore reference-free evaluation methods that prompt ChatGPT to score the quality of model-generated texts without a pre-defined ground truth. Chiang et al. (Chiang and Lee, 2023) show that LLM evaluation can be stable over prompt formatting and consistent with human experts. Jain et al. (Jain et al., 2023) propose self-supervised evaluation strategies to assess LLM properties without benchmarks or human annotations. Ye et al. propose FLASK (Ye et al., 2023), a fine-grained LLM evaluation protocol that decomposes coarse-level scoring to instance-wise skill assessment. FLASK prompts GPT-4 to assign a score with a rationale for each alignment skill of the target FM-generated text instance. From a data-centric perspective, the benchmark FM with extensive knowledge and human alignment embedded in its massive pre-training data can evaluate the performance of other FMs.
@@ 9.3. Bias
From a model perspective, healthcare stakeholders and developers should acknowledge the inherent bias present in FM architectures. Currently, FM architectures and training patterns lack defense and detection capabilities against adversarial manipulation. FMs could propagate all knowledge learned from training data without a proper mechanism to identify and address biased knowledge. This vulnerability can potentially hurt healthcare stakeholders by generating false or harmful information. Meanwhile, adversarial attack training is promising to enhance the harmful information defense and detection capability (Ferrara, 2023). Furthermore, the bias of human-AI alignment remains in FM development, where matching human value and model output will be a long-term challenge. To mitigate this bias, strong guidance from human experts in dataset curation and FM outcome evaluation is necessary in the healthcare domain. Human experts can provide reliable feedback to FM developers for identifying and addressing bias-related issues (Ferrara, 2023).